# Towards Depth Foundation Model: Recent Trends in Vision-Based Depth Estimation

### v1

@@
Abstract
Depth estimation is a fundamental task in 3D computer vision, crucial for applications such as 3D reconstruction, free-viewpoint rendering, robotics, autonomous driving, and AR/VR technologies. Traditional methods relying on hardware sensors like LiDAR are often limited by high costs, low resolution, and environmental sensitivity, limiting their applicability in real-world scenarios. Recent advances in vision-based methods offer a promising alternative, yet they face challenges in generalization and stability due to either the low-capacity model architectures or the reliance on domain-specific and small-scale datasets. The emergence of scaling laws and foundation models in other domains has inspired the development of ”depth foundation models”—deep neural networks trained on large datasets with strong zero-shot generalization capabilities. This paper surveys the evolution of deep learning architectures and paradigms for depth estimation across the monocular, stereo, multi-view, and monocular video settings. We explore the potential of these models to address existing challenges and provide a comprehensive overview of large-scale datasets that can facilitate their development. By identifying key architectures and training strategies, we aim to highlight the path towards robust depth foundation models, offering insights into their future research and applications.
Index Terms:
Depth Estimation, Foundation Models, 3D Computer Vision
@@ 1 Introduction
Figure 1: Scaling trends in model capacity and data volume for depth estimation. Each point represents a published method, positioned by its approximate model size (bottom axis, logarithmic scale of parameter count) and dataset size (left axis, logarithmic scale of training images), and colored by publication year. Early models (lightest shading, before 2020) relied on sub-million-parameter networks trained on only thousands of images, yielding limited generalization. Over 2020–2021 (mid-tone), methods grew to tens of millions of parameters and hundred-thousand-image datasets. The most recent models (darkest shading, 2022–2025) scale further into the billions of parameters and multi-million to ten-million-image with strong generalization ability and demonstrating the potential evolution toward depth foundation models.
@@
Domain Tasks Dataset Name Dynamic Scenes# Frames# Resolution Anno
Mono Stereo MV Video
Real-World A2D2 [1]   Dynamic 3 394K $1920\times 1208$ Sparse
Arogoverse2 [2, 3]   Dynamic 1000 2.14M $2048\times 1550$ Sparse
CityScapes [4]   Dynamic 2975 89K $2048\times 1024$ Stereo
DDAD [5]   Dynamic 200 98K $1936\times 1216$ Sparse
DIML [6]    Static 200+ 2M $1334\times 756$ Sparse
DIODE [7]    Static 30 27K $1024\times 768$ Sparse
DSEC [8]   Dynamic 53 26K $1440\times 1080$ Sparse
HM3D [9]   Static 1000 Dense
iBims-1 [10]   Static 20 54 $640\times 480$ Sparse
Lyft [11]   Dynamic 366 158K $1224\times 1024$ Sparse
Mapillary PSD [12]   Dynamic 50K 750K $1920\times 1080$ Sparse
NuScenes [13]   Dynamic 1K 40K $1600\times 900$ Sparse
NYU [14]   Static 464 435K $640\times 480$ Dense
Pandaset [15]   Dynamic 103 8K $1920\times 1080$ Sparse
Replica[16]   Static 18 36K $1200\times 680$ Dense
Taskonomy[17]   Static 600 4.5M $512\times 512$ Dense
KITTI [18]   Dynamic 22 41K $1242\times 375$ Sparse
DrivingStereo [19]   Dynamic 184 180K $1762\times 800$ Sparse
InStereo2K [20]   Static 50 2K $1080\times 860$ Dense
Argoverse [21]   Dynamic 113 6,624 $2056\times 2464$ Sparse
ETH3D [22]    Static 25 1K $6233\times 4146$ Dense
Waymo [23]   Dynamic 1,150 160K $1920\times 1280$ Sparse
UASOL [24]   Static 676 160.9K $2280\times 1282$ Dense
DTU [25]   Static 124 42.5K $1600\times 1200$ Dense
BlendedMVS [26]     Static 113 17k $2048\times 1536$ Dense
Tanks and Temples [27]    Static 21 147K $3840\times 2160$ Sparse
WildRGBD [28]   Static 23K 6M $480\times 640$ Dense
MVImgNet [29]   Static 220K 6.8M $1080\times 1920$ Dense
ARKitScenes [30]   Static 5,047 450M $256\times 192$ Dense
ARKitScenes-HighRes [30]   Static 5,047 450M $1920\times 1440$ Dense
Matterport3D [31]   Static 90 194.4K $1280\times 1024$ Dense
ScanNet [32]   Static 1,513 2.5M $640\times 480$ Dense
ScanNet++ [33]   Static 1,858 3.7M+ $1920\times 1440$ Dense
Synthetic Sintel [34]    Dynamic 10 1K $1024\times 436$ Dense
SceneFlow [35]     Dynamic 9 40K $960\times 540$ Dense
CREStereo [36]   Static 0 103K $1920\times 1080$ Dense
FallingThings [37]    Dynamic 3 62K $960\times 540$ Dense
FSD [38]     Dynamic 12 1M $1280\times 720$ Dense
UnrealStereo4K [39]    Static 8 7,720 $3840\times 2160$ Dense
Spring [40] Dynamic 47 6K $1920\times 1080$ Dense
TartanAir [41]    Static 163 1M $640\times 480$ Dense
VirtualKITTI2 [42]   Dynamic 5 25K $1242\times 375$ Dense
MatrixCity [43]   Static 3K 519K $1000\times 1000$ Dense
MVS-Synth [44]   Dynamic 120 12K $1920\times 1080$ Dense
3D Ken Burns [45]   Static 32 536K $512\times 512$ Dense
Dynamic Replica [46]   Dynamic 484 145K $1280\times 720$ Dense
OmniObject3D [47]   Static 6K 600K $800\times 800$ Dense
IRS [48]   Static 70 100K $960\times 540$ Dense
PointOdyssey [49]    Dynamic 131 200K $540\times 960$ Dense
BEDLAM [50]    Dynamic 10,450 380K $1280\times 720$ Dense
TABLE I: Datasets. The term   indicates that the scenes in these datasets are object-centric. Similarly,   and   refer to indoor and outdoor scenes, respectively. The depth annotations can be classified as dense or sparse, depending on whether most pixels have corresponding depth values. †: HM3D is a digital twin dataset created from real-world data, with the number of frames and resolutions being user-defined.
@@ 2.1 Definition of Depth Foundation Models
We provide a brief overview of the development of foundation models in language model to facility the understanding of depth foundation models. The field of language model have experienced explosive growth with the establishment of foundation models in recent years. This progress stems from the ability of these models to learn universal language and patterns from massive datasets, enabling them to generalize powerfully across various downstream tasks.
Convolutional neural networks and long short-term memory networks [51] plays as the main role at the early stage of language models, with limited network and data scales. The concept of Word Embeddings [52] and the introduction of the self-attention Mechanism [53], allowed the model to process all words in a sequence simultaneously, vastly improving parallel computation efficiency and the ability to capture long-range dependencies. The original Transformer model had a relatively small number of parameters, but its architecture laid the groundwork for subsequent large-scale models. BERTs [54] and GPTs [55] can be considered as the beginning of foundation models in large language models (LLMs). Proposed by Google, BERT is a bidirectional pre-trained model based on the Transformer architecture, enabling better understanding the polysemy of words in a sentence. Bert is trained on Toronto BookCorpus (800 million words) and English Wikipedia (2.5 billion words). BERT-Base has 110 million parameters, and BERT-Large has 340 million parameters. Proposed by OpenAI, GPT is a unidirectional generative pre-trained model based on the Transformer architecture. GPT models learn language patterns by predicting the next word, excelling in text generation tasks. The GPT-3 is trained on a dataset which is larger than 45 TB, along with 175 billion parameters.
@@ 3 Overview of Depth Estimation
In this section, we provide a overlook of paradigms and datasets used in depth models.
@@
Datasets.  We summarize the datasets commonly used in depth estimation tasks in Tab. I. These datasets can be categorized into two classes: real-world captured and synthesized. Each dataset may be applicable to multiple tasks, and we specify the tasks associated with each dataset in the table. Furthermore, we present details about the scenes within the datasets, including whether they provide metric information, if they include dynamic scenes, the number of scenes and frames, the image resolution, and whether the annotations are dense, based on the proportion of pixels that have corresponding depth values.
@@ 4 Monocular Image Depth Estimation
Evolution of method paradigms.  Parallel advancements emerged in depth representation and learning strategies. Early methods [60, 61, 62] focused on absolute depth prediction but faced scale ambiguity challenges, prompting innovations like scale-invariant loss [67, 68] for relative depth estimation. Another advancement came through bins classification approaches [69, 70, 71, 72, 73, 74], which discretize the continuous depth space into multiple bins. This reformulates depth estimation as a per-pixel classification problem, effectively handling non-uniform depth distributions and capturing uncertainty in predictions, particularly at depth discontinuities. Recent breakthroughs [75, 76, 77, 78, 79] integrate camera parameters to resolve metric ambiguity while maintaining scale awareness. The data paradigm evolved through large-scale multi-view stereo (MVS) datasets like MegaDepth [80], enabling unprecedented generalization. Supervision methods expanded beyond fully labeled training: self-supervised approaches exploited photometric consistency in stereo sequences [81], while Depth Anything [82] demonstrated pseudo-label distillation’s effectiveness for knowledge transfer. Multi-task frameworks [83, 84, 85, 86] further enhanced robustness through joint depth-flow-pose estimation with geometric constraints, complemented by geometry-aware losses like Virtual Normal that explicitly enforce surface regularity. This multifaceted progression underscores how architectural innovation, representation learning, and supervision paradigms collectively advance monocular depth estimation toward human-level scene understanding.
@@
Evolution towards foundation models.  Models for monocular depth estimation are becoming larger and more data-intensive, evolving towards foundation models. In terms of Model Size, advanced foundation models primarily utilize two key methods: Vision Transformer (ViT) architectures [87] and diffusion-based generative models [88]. Dense Prediction Transformers (DPT) are the main architecture for modern depth estimation [63], featuring 343 million parameters. Unlike older fully convolutional networks, DPT’s ViT backbone keeps high-resolution features and a global view throughout, leading to more detailed and consistent depth estimates. Pre-Trained Diffusion models are used as strong depth estimators to improve how well monocular depth estimation works, thanks to their existing visual knowledge, which ranges from 200 million to 1 billion parameters. For example, Marigold [66] slightly modifies text-to-image latent diffusion models (like Stable Diffusion v2 [89]) to predict depth from images. In terms of Data scale, powerful hardware and better cameras have led to many high-quality depth datasets, typically ranging from thousands to millions of images; further details can be found in Tab. I. The current trend is to train models on large datasets to make them more adaptable. When training for large-scale depth estimation, models typically use loss functions that ignore scale and shift differences in the data [67]. This means models learn to predict relative depth, making relative depth estimation a standard task in computer vision. However, metric depth estimation, which provides absolute distance measurements, is crucial for real-world applications. Currently, there are two main ways to get metric depth: 1) Methods like those in [76, 77, 79, 78, 90, 91] combine camera intrinsic estimations with relative depth predictions to get metric-scaled outputs through geometric calculations. 2) Directly learning metric depth: Approaches such as those in [92, 93, 94] train models directly on data that includes scale information, allowing them to predict metric depth without extra steps.

### v2

@@
[…]
The first two authors contributed equally to this work.
$1\qquad$ Zhejiang University, China.
$2\qquad$ Shanghai AI Lab, China.
$3\qquad$ Shenzhen University, China.
Authors’ emails: zhenx@zju.edu.cn, hongy_zhou@zju.edu.cn, pengsida@zju.edu.cn, haotongl@outlook.com, guohaoyu@pjlab.org.cn, jhshao@zju.edu.cn, yangps0306@gmail.com, yangqinglin@zju.edu.cn, 12231099@zju.edu.cn, hexingyi8@gmail.com, accwyf@gmail.com, ywang24@zju.edu.cn, ruizhen.hu@szu.edu.cn, yiyi.liao@zju.edu.cn, xwzhou@zju.edu.cn, bao@cad.zju.edu.cn
Manuscript received: 2025-01-01; accepted: 2025-01-01
@@ 1 Introduction
Figure 1: Scaling trends in model capacity and data volume for depth estimation. Each point represents a published method, positioned by its approximate model size (bottom axis, logarithmic scale of parameter count) and dataset size (left axis, logarithmic scale of training images), and colored by publication year. Early models (lightest shading, before 2020) relied on sub-million-parameter networks trained on only thousands of images, yielding limited generalization. Over 2020–2021 (light blue shading), methods grew to tens of millions of parameters and hundreds of thousands of image datasets. In 2022-2023 (gray shading), methods continued to increase in size and scalable training, with around one billion parameters and datasets ranging from several million to ten million images. The most recent models (darkest shading, 2024–2025) further scale up to the billions of parameters and utilize over ten million images, with strong generalization ability, and demonstrate the potential evolution toward depth foundation models.
Domain Tasks Dataset Name Dynamic Metric Scenes# Frames# Resolution Anno License
Mono Stereo MV Video
Real-World
A2D2 [1]   Dynamic Yes 3 394K $1920\times 1208$ Sparse CC BY-ND 4.0
Arogoverse2 [2, 3]   Dynamic Yes 1000 2.14M $2048\times 1550$ Sparse CC BY-NC-SA 4.0
CityScapes [4]   Dynamic Yes 2975 89K $2048\times 1024$ Sparse Non-Commercial
DDAD [5]   Dynamic Yes 200 98K $1936\times 1216$ Sparse CC BY-NC-SA 4.0
DIML [6]    Static Yes 200+ 2M $1334\times 756$ Sparse Non-Commercial
DIODE [7]    Static Yes 30 27K $1024\times 768$ Sparse MIT
DSEC [8]   Dynamic Yes 53 26K $1440\times 1080$ Sparse CC BY-NC-SA 4.0
HM3D [9]   Static Yes 1000 Dense MIT
iBims-1 [10]   Static Yes 20 54 $640\times 480$ Sparse CC BY 4.0
Lyft [11]   Dynamic Yes 366 158K $1224\times 1024$ Sparse CC0 1.0
Mapillary PSD [12]   Dynamic Yes 50K 750K $1920\times 1080$ Sparse CC BY-NC-SA
NuScenes [13]   Dynamic Yes 1K 40K $1600\times 900$ Sparse Non-Commercial
NYU [14]   Static Yes 464 435K $640\times 480$ Dense GPL
Pandaset [15]   Dynamic Yes 103 8K $1920\times 1080$ Sparse CC0
Taskonomy[16]   Static Yes 600 4.5M $512\times 512$ Dense MIT
KITTI [17]   Dynamic Yes 22 41K $1242\times 375$ Sparse CC BY-NC-SA 3.0
DrivingStereo [18]   Dynamic Yes 184 180K $1762\times 800$ Sparse MIT
InStereo2K [19]   Static Yes 50 2K $1080\times 860$ Dense Non-Commercial
Argoverse [20]   Dynamic Yes 113 6,624 $2056\times 2464$ Sparse CC BY-NC-SA 4.0
ETH3D [21]    Static Yes 25 1K $6233\times 4146$ Dense CC BY-NC-SA 4.0
Waymo [22]   Dynamic Yes 1,150 160K $1920\times 1280$ Sparse Non-Commercial
UASOL [23]   Static Yes 676 160.9K $2280\times 1282$ Dense CC BY-NC-SA 3.0
DTU [24]   Static Yes 124 42.5K $1600\times 1200$ Dense MIT
BlendedMVS [25]     Static No 113 17k $2048\times 1536$ Dense CC BY 4.0
WildRGBD [26]   Static Yes 23K 6M $480\times 640$ Dense MIT
MVImgNet [27]   Static No 220K 6.8M $1080\times 1920$ Dense CC BY-NC 4.0
ARKitScenes [28]   Static Yes 5,047 450M $256\times 192$ Dense Non-Commercial
ARKitScenes-HighRes [28]   Static Yes 5,047 450M $1920\times 1440$ Dense Non-Commercial
Matterport3D [29]   Static Yes 90 194.4K $1280\times 1024$ Dense MIT
Replica[30]   Static Yes 18 36K $1200\times 680$ Dense Non-Commercial
Dynamic Replica [31]   Dynamic Yes 484 145K $1280\times 720$ Dense CC BY-NC 4.0
ScanNet [32]   Static Yes 1,513 2.5M $640\times 480$ Dense Non-Commercial
ScanNet++ [33]   Static Yes 1,858 3.7M+ $1920\times 1440$ Dense Non-Commercial
Synthetic
Sintel [34]    Dynamic No 10 1K $1024\times 436$ Dense CC BY 3.0
SceneFlow [35]     Dynamic No 9 40K $960\times 540$ Dense CC BY 4.0
CREStereo [36]   Static No 0 103K $1920\times 1080$ Dense Apache-2.0
FallingThings [37]    Dynamic No 3 62K $960\times 540$ Dense CC BY-NC-SA 4.0
FSD [38]     Dynamic No 12 1M $1280\times 720$ Dense Non-Commercial
UnrealStereo4K [39]    Static No 8 7,720 $3840\times 2160$ Dense MIT
Spring [40] Dynamic No 47 6K $1920\times 1080$ Dense CC BY 4.0
TartanAir [41]    Static Yes 163 1M $640\times 480$ Dense CC BY 4.0
VirtualKITTI2 [42]   Dynamic Yes 5 25K $1242\times 375$ Dense CC BY-NC-SA 4.0
MatrixCity [43]   Static Yes 3K 519K $1000\times 1000$ Dense Apache-2.0
MVS-Synth [44]   Dynamic No 120 12K $1920\times 1080$ Dense Non-Commercial
3D Ken Burns [45]   Static No 32 536K $512\times 512$ Dense CC BY-NC-SA 4.0
OmniObject3D [46]   Static Yes 6K 600K $800\times 800$ Dense CC BY 4.0
IRS [47]   Static Yes 70 100K $960\times 540$ Dense Apache 2.0
PointOdyssey [48]    Dynamic No 131 200K $540\times 960$ Dense MIT
BEDLAM [49]    Dynamic Yes 10,450 380K $1280\times 720$ Dense Non-Commercial
Table 1: Datasets. The term   indicates that the scenes in these datasets are object-centric. Similarly,   and   refer to indoor and outdoor scenes, respectively. The depth annotations can be classified as dense or sparse, depending on whether most pixels have corresponding depth values. †: HM3D is a digital twin dataset created from real-world data, with the number of frames and resolutions being user-defined.
@@
Task
Method Year Model Type Key Paradigm Model Size Data Size Benchmarks
NYUv2 [14] (Rel.) $\downarrow$ KITTI [17, 17] (Rel.) $\downarrow$
Monocular
Multi-Scale Net [51] 2014 CNN Direct Regression $\sim$70M $\sim$150K 0.190
LRC [52] 2016 CNN Direct Regression $\sim$30M $\sim$50K 0.114
DORN [53] 2018 CNN Depth Classification $\sim$50M $\sim$200K 0.115 0.072
PAD-Net [54] 2018 CNN Multiple Tasks $\sim$25M $\sim$3K 0.120
MiDaS [55] 2019 CNN Affine Invariant $\sim$90M $\sim$2M 0.111 0.236
LeReS [56] 2020 CNN Canonical Camera $\sim$90M $\sim$350K 0.090 0.149
DPT [57] 2021 ViT Scaling Law $\sim$350M $\sim$1.5M 0.098 0.100
Marigold [58] 2024 Diffusion Scaling Law $\sim$800M $\sim$75K 0.055 0.099
DA [59] 2024 ViT Scaling Law $\sim$600M $\sim$63.5M 0.0382 0.0504
DAV2 [60] 2024 ViT Scaling Law $\sim$2B $\sim$62.5M 0.0416 0.0677
MoGE [61] 2024 ViT Scaling Law $\sim$350M $\sim$4M 0.0292 0.0394
UniDepth [62] 2024 ViT Metric Depth $\sim$350M $\sim$4M 0.0340/0.116 0.0350/0.0469
UniDepthV2 [63] 2025 ViT Metric Depth $\sim$350M $\sim$4M 0.0296/0.106 0.0385/0.0858
MoGE2 [64] 2025 ViT Metric Depth $\sim$350M $\sim$4M 0.0289/0.073 0.0375/0.1810
KITTI 2015 [17] (D1) $\downarrow$
Stereo
GCNet [65] 2017 CNN Cost Volume $\sim$4M $\sim$35K 2.87%
STTR [66] 2020 ViT Attention $\sim$3M $\sim$22K 2.01%
LEAStereo [67] 2020 CNN Attention $\sim$2M $\sim$20K 1.65%
RAFT-Stereo [68] 2021 CNN+GRU Iterative Optim. $\sim$10M $\sim$35K 1.96%
Mono2Stereo [69] 2024 CNN Scaling Law $\sim$5M $\sim$600K 1.58%
DTU [70] (Overall Err.) $\downarrow$
Multi-view
Gipuma [71] 2015 Handcrafted Handcrafted Algo. $\sim$0 0 0.283
MVSNet [72] 2018 CNN Cost Volume $\sim$4M $\sim$27K 0.462
DUSt3R [73] 2023 ViT Scaling Law $\sim$400M $\sim$17M 1.741
MVSFormer++ [74] 2024 ViT Scaling Law $\sim$40M $\sim$40K 0.281
MASt3R [75] 2024 ViT Scalable Law $\sim$400M $\sim$20M 0.374
VGGt [76] 2025 ViT Scaling Law $\sim$1B $\sim$30M 0.382
ScanNet [32] (Rel.) $\downarrow$ KITTI [17, 17] (Rel.) $\downarrow$
Video
CVD [77] 2020 CNN Test-Time Optim. $\sim$25M $\sim$120K 0.073
MAMo [78] 2023 ViT Temporal Corr. $\sim$200M $\sim$60K 0.049
NVDS [79] 2023 ViT Scaling Law $\sim$100M $\sim$2.2M 0.187 0.253
ChronoDepth [80] 2024 Diffusion Scaling Law $\sim$2B $\sim$240K 0.159 0.151
DepthCrafter [81] 2024 Diffusion Scaling Law $\sim$2B $\sim$25M 0.125 0.110
Buffer Anytime [82] 2024 ViT+Diffusion Scaling Law $\sim$300M $\sim$26M 0.123 0.119
Table 2: Key methods in depth estimation. Model size and data size are approximate. Benchmark results are collected from the corresponding papers. Sections marked with different background colors (  light gray or   medium gray) indicate different evaluation protocols used by the papers.
@@ 2.1 Definition of Depth Foundation Models
We provide a brief overview of the development of foundation models in language models and self-supervised vision-based tasks to facilitate the understanding of the depth of foundation models. The field of language models has experienced explosive growth with the establishment of foundation models in recent years. This progress stems from the ability of these models to learn universal language and patterns from massive datasets, enabling them to generalize powerfully across various downstream tasks. Moreover, development on scalable architectures and self-supervision tasks has also enabled the emergence of vision foundation models, which are trained on a vast amount of image or video data to facilitate the 2D or 3D perception task.
Convolutional neural networks and long short-term memory networks [83] played a main role at the early stage of language models, with limited network and data scales. The concept of Word Embeddings [84] and the introduction of the self-attention Mechanism [85] allowed the model to process all words in a sequence simultaneously, vastly improving parallel computation efficiency and the ability to capture long-range dependencies. The original Transformer model had a relatively small number of parameters, but its architecture laid the groundwork for subsequent large-scale models. BERTs [86] and GPTs [87] can be considered as the beginning of foundation models in large language models (LLMs). Proposed by Google, BERT is a bidirectional pre-trained model based on the Transformer architecture, enabling better understanding of the polysemy of words in a sentence. Bert is trained on Toronto BookCorpus (800 million words) and English Wikipedia (2.5 billion words). BERT-Base has 110 million parameters, and BERT-Large has 340 million parameters. Proposed by OpenAI, GPT is a unidirectional generative pre-trained model based on the Transformer architecture. GPT models learn language patterns by predicting the next word, excelling in text generation tasks. The GPT-3 is trained on a dataset that is larger than 45 TB, along with 175 billion parameters.
Vision foundation models have recently emerged as powerful tools for a wide range of visual perception tasks, closely following the development of scalable model architectures in language models. These models, such as DINO [88, 89, 90], MAE [91], and SAM [92, 93], are typically trained on massive image or video datasets using self-supervised or weakly supervised objectives. By leveraging scalable architectures (e.g., Vision Transformers) and large-scale pretraining, vision foundation models learn rich and generalizable visual representations that can be transferred to downstream tasks, including classification, detection, segmentation, and even 3D perception. Notably, recent diffusion-based vision models [94, 95] have also demonstrated strong capabilities in generative and dense prediction tasks. These advances in vision foundation models provide the architectural and data-driven basis for the development of depth foundation models, enabling improved generalization and scalability for depth estimation across diverse domains.
@@ 2.2 Depth Estimation Tasks
Moreover, we additionally define the task of metric depth estimation as one that requires the model to not only output the relative depth map, but also predict its real-world scale and offset. For the tasks of stereo, multi-view, and monocular video depth estimation, depth estimation requires predicting depth maps that, when unprojected to the world coordinate system, align across different views. When the real-world metric-scale camera parameters are known, these three tasks automatically become metric depth estimation tasks. Thus, we only distinguish between relative and metric depth estimation in the monocular depth estimation task.
@@ 3 Overview of Depth Estimation
In this section, we provide an overlook of paradigms and datasets used in depth models.
@@
Datasets.  We summarize the datasets commonly used in depth estimation tasks in table 1. These datasets can be categorized into two classes: real-world captured and synthesized. Each dataset may be applicable to multiple tasks, and we specify the tasks associated with each dataset in the table. Furthermore, we present details about the scenes within the datasets, including whether they provide metric information, if they include dynamic scenes, the number of scenes and frames, the image resolution, and whether the annotations are dense, based on the proportion of pixels that have corresponding depth values.
@@ 4 Monocular Image Depth Estimation
Evolution of method paradigms.  Parallel advancements emerged in depth representation and learning strategies. Early methods [51, 100, 101] focused on absolute depth prediction but faced scale ambiguity challenges, prompting innovations like scale-invariant loss [55, 104] for relative depth estimation. Another advancement came through bin classification approaches [53, 105, 106, 107, 108, 109], which discretize the continuous depth space into multiple bins. This reformulates depth estimation as a per-pixel classification problem, effectively handling non-uniform depth distributions and capturing uncertainty in predictions, particularly at depth discontinuities. Recent breakthroughs [56, 110, 111, 62, 112] integrate camera parameters to resolve metric ambiguity while maintaining scale awareness. The data paradigm evolved through large-scale multi-view stereo (MVS) datasets like MegaDepth [113], enabling unprecedented generalization. Supervision methods expanded beyond fully labeled training: self-supervised approaches exploited photometric consistency in stereo sequences [114], while Depth Anything [59] demonstrated pseudo-label distillation’s effectiveness for knowledge transfer. Multi-task frameworks [115, 54, 116, 117] further enhanced robustness through joint depth-flow-pose estimation with geometric constraints, complemented by geometry-aware losses like Virtual Normal that explicitly enforce surface regularity. This multifaceted progression underscores how architectural innovation, representation learning, and supervision paradigms collectively advance monocular depth estimation toward human-level scene understanding.
@@
Evolution towards foundation models.  Models for monocular depth estimation are becoming larger and more data-intensive, evolving towards foundation models. In terms of Model Size, advanced foundation models primarily utilize two key methods: Vision Transformer (ViT) architectures [118] and diffusion-based generative models [95]. Dense Prediction Transformers (DPT) are the main architecture for modern depth estimation [57], featuring 343 million parameters. Unlike older fully convolutional networks, DPT’s ViT backbone keeps high-resolution features and a global view throughout, leading to more detailed and consistent depth estimates. Pre-Trained Diffusion models are used as strong depth estimators to improve how well monocular depth estimation works [119, 58, 120, 121], thanks to their existing visual knowledge, which ranges from 200 million to 1 billion parameters. For example, Marigold [58] modifies text-to-image latent diffusion models (like Stable Diffusion v2 [94]) to predict depth from images. In terms of Data scale, powerful hardware and better cameras have led to many high-quality depth datasets, typically ranging from thousands to millions of images; further details can be found in table 1. The current trend is to train models on large datasets to make them more adaptable. When training for large-scale depth estimation, models typically use loss functions that ignore scale and shift differences in the data [55]. This means models learn to predict relative depth, making relative depth estimation a standard task in computer vision. However, metric depth estimation, which provides absolute distance measurements, is crucial for real-world applications. Currently, there are two main ways to get metric depth: 1) Methods like those in [110, 111, 112, 62, 61, 64, 122] combine camera intrinsic estimations with relative depth predictions to get metric-scaled outputs through geometric calculations. 2) Directly learning metric depth: Approaches such as those in [123, 124, 125] train models directly on data that includes scale information, allowing them to predict metric depth without extra steps.